# LMME3DHF: Benchmarking and Evaluating Multimodal 3D Human Face Generation with LMMs

### v2

@@ 5.3. Performance on Distortion-aware Saliency Prediction
To evaluate performance on the distortion-aware saliency prediction task, we compare LMME3DHF with current state-of-the-art saliency prediction models, including both traditional and deep learning-based approaches. As shown in Table 3, LMME3DHF significantly outperforms all baseline models across various evaluation metrics. This superior performance can be attributed to the nature of our task, which differs from conventional saliency prediction. Unlike general saliency detection that focuses on broad visual attention, our task targets sparse, distortion-aware saliency, identifying precise regions associated with visual distortions. As a result, commonly used saliency prediction models are not well-suited for this specialized task, leading to limited performance in comparison. The visual comparisons presented in Figure 8 and the predictions on sample images from the Gen3DHF dataset in Figure 9 further emphasize that LMME3DHF outperforms other saliency prediction models, clearly demonstrating its superior ability to accurately localize distortion-aware salient regions.
@@ 5.4. Ablation Study
We conduct ablation experiments to validate the contributions of the key components in our proposed LMME3DHF framework. Table 4 shows the results for saliency-distortion-aware decoder. Experiments (1) show that removing features from the language decoder causes a noticeable drop in performance. Experiment (2) further emphasizes this scenario by setting an empty prompt, thereby demonstrating the critical role of textual guidance. Experiment (3) shows that using only fused multimodal features enhances the model’s ability to predict distortion-aware saliency. Experiment (4) our approach that combines both visual and multimodal features and achieves the highest performance, confirming the effectiveness of integrating both feature types. Results for score prediction and visual question answering are summarized in Table 5. Experiment (1) which fine-tuning only the vision encoder, yields the weakest performance across both tasks. Experiment (2) demonstrates that fine-tuning the LLM significantly improves performance, with the improvement being especially pronounced in the visual question answering task, which is inherently text-based. Experiment (3) shows a significant improvement in score prediction tasks, highlighting the effectiveness of the quality regression module. Experiment (4) performs comparably to Experiment (2), suggesting that the vision encoder is less critical in improving evaluation results. Finally, Experiments (5) and (6) achieve the best overall performance. Among them, we select the configuration from Experiment (5) as the default setup for LMME3DHF, as it balances strong performance with computational efficiency.

### v3

@@ 5.3. Performance on Distortion-aware Saliency Prediction
To evaluate performance on the distortion-aware saliency prediction task, we compare LMME3DHF with current state-of-the-art saliency prediction models, including both traditional and deep learning-based approaches. As shown in Table 3, LMME3DHF significantly outperforms all baseline models across various evaluation metrics. Unlike general saliency detection that focuses on broad visual attention, our task targets sparse, distortion-aware saliency, identifying precise regions associated with visual distortions. As a result, commonly used saliency prediction models are not well-suited for this specialized task, leading to limited performance in comparison. The visual comparisons presented in Figure 8 and the predictions on sample images from the Gen3DHF dataset in Figure 9 further emphasize that LMME3DHF outperforms other saliency prediction models, clearly demonstrating its superior ability to accurately localize distortion-aware salient regions.
@@ 5.4. Ablation Study
We conduct ablation experiments to validate the contributions of the key components in our proposed LMME3DHF framework. Table 4 shows the results for saliency-distortion-aware decoder. Experiments (1) show that removing features from the language decoder causes a noticeable drop in performance. Experiment (2) further emphasizes this scenario by setting an empty prompt, thereby demonstrating the critical role of textual guidance. Experiment (3) shows that using only fused multimodal features enhances the model’s ability to predict distortion-aware saliency. Experiment (4) our approach that combines both visual and multimodal features achieves the highest performance, confirming the effectiveness of integrating both feature types. Results for score prediction and visual question answering are summarized in Table 5. Experiment (1) which fine-tuning only the vision encoder, yields the weakest performance across both tasks. Experiment (2) demonstrates that fine-tuning the LLM significantly improves performance, with the improvement being especially pronounced in the visual question answering task, which is inherently text-based. Experiment (3) shows a significant improvement in score prediction tasks, highlighting the effectiveness of the quality regression module. Experiment (4) performs comparably to Experiment (2), suggesting that the vision encoder is less critical in improving evaluation results. Finally, Experiments (5) and (6) achieve the best overall performance. Among them, we select the configuration from Experiment (5) for LMME3DHF, as it balances strong performance with computational efficiency.
@@ 6. CONCLUSION
acknowledgements
This work was supported in part by the National Natural Science Foundation of China under Grant 62271312 and Grant 62132006, and in part by STCSM under Grant 22DZ2229005.